Find the equation of tangent and normal of curve x^2/3+y^2/3=a^2/3 at (h,k)

Grade:12th pass

3 Answers

Arun
4 years ago
Dear student
 
To find the tangents to a give curve,differentiate the the give curve....
On differentiating...we get slope of the tangent..
 
Differentiating with respect to x,
2/3(x-1/3) + 2/3(y-1/3) dy/dx = 0
 
dy/dx = -x-1/3/y-1/3
NOW FIND slope at (h,k) and then write equation of tangent
Vikas TU
4 years ago
Dear student 
Please follow this link 
Good Luck 
Aditya Gupta
4 years ago
the correct answer is 
replace x^2 by xh, and y^2 by yk
so, eqn of tangent is 
 xh/3+yk/3=a^2/3 (slope = – h/k)
normal can be written as
y – k= m(x – h)
where m*( – h/k)= – 1
or m= k/h
or yh – kh= kx – kh
or y= kx/h
kindly approve :))